choose the equation of the line passing through the point (4,5) and perpendicular to y=1/4x+9
Answer by mananth(16946)

- 1/ 4 x + 1 y = 9
Find the slope of this line

1 y = 1/ 4 x + 9

Compare this equation with y=mx+b,m= slope & b= y intercept
slope m = 1/4

The slope of a line perpendicular to the above line will be the negative reciprocal -4
Because m1*m2 =-1
The slope of the required line will be -4

m= -4 ,point ( 4 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = -16 + b
b= 21
m= -4
The required equation is y = -4 x + 21
m.ananth@hotmail.ca